The speed of a passenger train is 16 mph faster than the speed of a freight train. The passennger train travels 300 miles in the same time it takes the freight train to travel 220 miles. What is the speed of each train?
Answer by mananth(16946)

freight train x 220 miles
Passenger train x+ 16 300 miles

t=d/r
220/x=300/(x+16)
Cross multiply
220(x+16)=300*x
220 x + 3520 = 300 x
300 x -220 x = 3520
80 x = 3520
/ 80
x = 44 freight train mph
60 Passenger train mph